A work-standards expert observes the amount of time required to prepare a sample of 10 business letters in an office with the following results listed in ascending order to the nearest minute: 5, 5, 5, 7, 9, 14, 15, 15, 16, 18. Determine the
Range
1
Expert's answer
2021-08-10T06:15:06-0400

(i) "n=10"


"5,5,5,7,9,14,15,15,16,18"

"Range=18-5=13"

(ii)


"mean =\dfrac{1}{15}(5+5+5+7+9+14+15+15"

"+16+18)=10.9"

(iii)

"median=\dfrac{9+14}{2}=11.5"

(iv)

"mode=5\ (3\ times)"

(v)


"Var(x)=\sigma^2=\dfrac{1}{10}((5-10.9)^2+(5-10.9)^2"

"+(5-10.9)^2+(7-10.9)^2+(9-10.9)^2"

"+(14-10.9)^2+(15-10.9)^2+(15-10.9)^2"

"+(16-10.9)^2+(18-10.9)^2=24.29"

(vi)


"\sigma=\sqrt{\sigma^2}=\sqrt{24.29}\approx4.9285"